OpenStudy (anonymous):

SOLVE 2a-6+5a= 10 for a

OpenStudy (m):

you want to separate variable a on on side 2a + 5a = 10 + 6 7a = 16 a = 16/7

OpenStudy (anonymous):

in decimal form?

OpenStudy (m):

2.285714...

OpenStudy (m):

use calculator on your computer for complete answer :)

OpenStudy (anonymous):

i got it i suck at math and i have no tutor......... im on homebound i do school work just at home

OpenStudy (m):

oh i see. this is a good place to get help for algebra :)

OpenStudy (m):

keep asking questions!

OpenStudy (anonymous):

i did...... im in geomatry so there are lots of shapes but i am asking algabra questions